# A Hausdorff-measure boundary element method for acoustic scattering by fractal screens

### v1

@@ 6 Numerical results
In this section we present numerical results that illustrate our methods and assess the sharpness of our theoretical predictions. The Hausdorff BEM (57) has been implemented in the Julia language [4] and the code is available at https://github.com/AndrewGibbs/IFSintegrals. In all examples described in this section, we consider the scattering of plane waves, i.e. the datum $g$ is as in (31) with $u^{i}(x)=e^{{\mathrm{i}}k\vartheta\cdot x}$ and $|\vartheta|=1$. We validate our implementation against a different method in §6.4, but in the rest of our experiments we use as a reference solution a more accurate Hausdorff-BEM solution with a large number of degrees of freedom $N_{\mathrm{ref}}$. Most of our experiments are for homogeneous attractors, in which case $N_{\mathrm{ref}}=M^{\ell_{\mathrm{ref}}}$ for some $\ell_{\mathrm{ref}}\in\mathbb{N}$. For Cantor sets $(M=2)$ we choose $\ell_{\mathrm{ref}}=15$, so that $N_{\mathrm{ref}}=32768$. For Cantor Dusts ($M=4$) we choose $\ell_{\mathrm{ref}}=8$, so that $N_{\mathrm{ref}}=65536$.
@@
Our implementation uses the quadrature rules described in §5.4. Precisely, we approximate the right-hand side in the linear system and linear functionals of the solution (the scattered field and far field) using the quadrature rules (99) and (100), respectively. We approximate the diagonal matrix elements $A_{ii}$ by (95)-(98). To approximate $A_{ij}$ with $i\neq j$ we use (94) with $h_{Q}$ replaced by $h_{Q,i,j}\geq h_{Q}$ given (in terms of $h_{Q}$) by (123)-(125). All these quadrature rules depend on the parameter $h_{Q}$. We choose $h_{Q}=C_{Q}h$, where the constant $C_{Q}$ is chosen to depend on the wavenumber as detailed below. While the requirement $h_{Q}=C_{Q}h$ is weaker (i.e. it requires fewer quadrature points) than the conditions $h_{Q}\leq C_{Q}h^{1+d/2}$ (see (113)) and $h_{Q}\leq C_{Q}h^{1+d}$ (see (111)) required by our theory, our numerical experiments suggest that, in practice, $h_{Q}=C_{Q}h$ is sufficient to achieve our theoretical convergence rates, even when using the reduced quadrature of Remark 5.19.
@@
We measure the accuracy of our BEM solutions in the $H^{-1/2}_{\Gamma}$ norm. These $H^{-1/2}_{\Gamma}$ norms are computed by expressing them in terms of a single-layer BIO with wavenumber $k={\mathrm{i}}$, which we denote $S^{\mathrm{i}}$, as in [16, Table 1]. Practically, we achieve this by assembling an approximation $A^{{\mathrm{i}},Q}$ to the Galerkin matrix $A^{\mathrm{i}}$ for the operator $S^{\mathrm{i}}$, analogous to (93), with $k={\mathrm{i}}$ and $N_{\mathrm{ref}}$ degrees of freedom, using quadrature approximations analogous to (95)-(98) and (94), choosing $h_{Q}$ the same as for the reference solution, with the reduced quadrature formulae (123)-(125) applied as if $k=1$. (While the quadrature convergence analysis in [25, §5] was presented only for $k>0$, one can check that the relevant results also extend, mutatis mutandis, to the case $k={\mathrm{i}}$.) Next, we view $\phi_{N}^{Q}$ as an element of the larger space $V_{N_{\mathrm{ref}}}$ and define $\vec{v}$} as the coefficient vector of $\phi_{N}^{Q}-\phi_{N_{\mathrm{ref}}}^{Q}$ in $V_{N_{\mathrm{ref}}}$. Then, arguing as in [16, Table 1] (and see Footnote 16), it follows that
@@ 6.1 $H^{-1/2}_{\Gamma}$-norm convergence, Cantor set ($n=1$)
In Figure 7 we show an example of the total and scattered fields for the middle-third case ($\rho=1/3$). In Figure 8 we present plots of the relative errors in the Hausdorff-BEM solution for $\ell=0,1,\ldots,14$, measured in $H^{-1/2}_{\Gamma}$ norm, using $\phi_{15}$ as reference solution, for various $\rho$ and $k$ values. In all cases the incident wave is the plane wave $u^{i}(x)={\mathrm{e}}^{{\mathrm{i}}k\vartheta\cdot x}$ with $\vartheta=(1,-1)/\sqrt{2}$. The experiments in Figures 7 and 8 are carried out using the quadrature rules described in §5.4, with the reduced quadrature of Remark 5.19, and with $h_{Q}=\rho^{2}h$ for $k=0.1$ and $h_{Q}=\rho^{4}h$ for $k=50$. This means that the off-diagonal entries of the BEM matrix require up to $16$ and $256$ evaluations of the Helmholtz fundamental solution for $k=0.1$ and $k=50$, respectively.
@@ 6.2 $H^{-1/2}_{\Gamma}$-norm convergence, Cantor dust ($n=2$)
a $d$-set with $d={\rm dim_{H}}(\Gamma)=\log{4}/\log{(1/\rho)}$, see Figure 1. Such a screen generates a non-zero scattered field if and only if $\rho>1/4$ (see [12, Example 8.2] and the discussion before Lemma 4.4). In Figure 9 we present results similar to those in Figure 8, albeit for a more restricted range of $\ell$ and a lower value of $k$ in the right-hand panel, due to the increased computational cost associated with the change from $M=2$ (Cantor set) to $M=4$ (Cantor dust). The Hausdorff-BEM solution $\phi_{\ell}$ corresponds to a mesh of $N=4^{\ell}$ elements of diameter $h=\sqrt{2}\rho^{\ell}$. The incident wave is the plane wave $u^{i}(x)={\mathrm{e}}^{{\mathrm{i}}k\vartheta\cdot x}$ with $\vartheta=(0,1,-1)/\sqrt{2}$. We use the quadrature rules described in §5.4, with the reduced quadrature of Remark 5.19, and with $h_{Q}=\rho^{2}h$ for both $k=0.1$ and $k=5$. As a result, the off-diagonal entries of the BEM matrix require up to $256$ evaluations of $\Phi$.
Figure 9: $H^{-1/2}_{\Gamma}$ relative errors for Hausdorff-BEM solutions on Cantor dusts. The legend inset in the right-hand panel applies to both panels.
@@
$\|\phi-\phi_{\ell}\|_{H^{-1/2}_{\Gamma}}=C\alpha^{\ell},\quad\ell\in\mathbb{N},$ (128)
for $\alpha:=(4\rho)^{-1/2}\in(1/\sqrt{2},1)$ and some $C>0$. Then, by the triangle inequality,
$C\alpha^{\ell}(1-\alpha^{\ell^{\prime}-\ell})\leq\|\phi_{\ell}-\phi_{\ell^{\prime}}\|_{H^{-1/2}_{\Gamma}}\leq C\alpha^{\ell}(1+\alpha^{\ell^{\prime}-\ell}),\qquad\ell,\ell^{\prime}\in\mathbb{N}_{0}\;\text{with}\;\ell<\ell^{\prime},$ (129)
@@
In Figure 11 we plot, as a better test of the sharpness of our theory, the norm $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}$, for $\ell=0,\ldots,7$, of the difference between Galerkin solutions at consecutive levels $\ell$. We observe, for each choice of $\rho$ and $k$, that, for large enough $\ell$ (how large depending on $\rho$ and $k$), $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}=c\alpha^{\ell}=c(4\rho)^{-\ell/2}$, for some constant $c>0$, also depending on $\rho$ and $k$, in agreement with (129) with $\ell^{\prime}=\ell+1$. Moreover, since $\|\phi-\phi_{\ell}\|_{H^{-1/2}_{\Gamma}}\to 0$ as $\ell\to\infty$ by Theorem 5.6, this observed convergence implies that
$\|\phi-\phi_{\ell}\|_{H^{-1/2}_{\Gamma}}\leq\sum_{m=\ell}^{\infty}\|\phi_{m}-\phi_{m+1}\|_{H^{-1/2}_{\Gamma}}=\frac{c}{1-\alpha}\,\alpha^{\ell},$
@@ 6.3 Convergence of the scattered field
In Figure 13 we show $L_{\infty}$ errors in the near- and far-field for scattering by Cantor sets and Cantor dusts, obtained by computing the maximum error over a suitable set of sample points. In more detail, define the parameter $\mathcal{N}(k):=10\max(k,2)$, which is always an even integer for the values of $k$ in our experiments. For the near-field, when $n=1$ we sample at $4\mathcal{N}$ points on the boundary of the square $(-1,2)\times(-1.5,1.5)$, and when $n=2$ we sample at $\mathcal{N}^{2}$ points on a uniform grid on the square $(-1,2)\times(-1,2)\times\{-1\}$ (recall that $\Gamma\subset[0,1]\times[0,1]\times\{0\}$). For the far-field, when $n=1$ we sample at $\mathcal{N}$ points on the circle $\mathbb{S}^{1}$, and when $n=2$ we sample at $\frac{\mathcal{N}}{2}\times\mathcal{N}$ points on the sphere $\mathbb{S}^{2}$, chosen such that the points form a uniform grid in spherical coordinate space $0,\pi]\times[0,2\pi]$.
@@ 6.4 Comparison against “prefractal BEM”
Figure 16: Ratio of errors of Hausdorff-BEM (${u}_{10}^{\infty}$) and prefractal-BEM ($\tilde{u}_{10}^{\infty}$) approximations of the far-field pattern $u^{\infty}$ with $\ell=10$ for scattering by Cantor sets with $k=0.1$, $k=30$ and $k=50$ and $\vartheta=(1/2,-\sqrt{3}/2)$, for a range of values of $\rho\in(0,1/2)$, measured in the $L_{\infty}$ (left panel) and $L_{2}$ (right panel) norms on $\mathbb{S}^{1}$. A value below 11 indicates that the Hausdorff BEM is more accurate, while a value above 11 indicates that the prefractal BEM is more accurate. In both cases $\ell_{\mathrm{ref}}=13$.
@@
In Figure 15 we show $L_{\infty}$ errors in the far-field pattern between the two methods for $\rho=0.1$, $1/3$ and $0.49$, with incident direction $\vartheta=(1/2,-\sqrt{3}/2)$ and wavenumbers $k=0.1$ and $k=50$ (with $h_{Q}=\rho^{6}h$ for $k=0.1$ and $h_{Q}=\rho^{8}h$ for $k=50$ for the Hausdorff BEM), obtained by computing the maximum error over 300 equally-spaced observation angles in $\mathbb{S}^{1}$. For each method we calculate the errors in two different ways, using, firstly, a Hausdorff-BEM reference solution, and, secondly, a prefractal-BEM reference solution; in both cases the reference solution is computed with $\ell_{\mathrm{ref}}=13$. In all cases, it is clear that both methods are converging to the same solution, validating both methods. Indeed, it appears that both methods converge at the same $2^{-\ell}$ rate. However, the results suggest that the relative accuracy of the two methods is dependent on the value of $\rho$, with the two methods having essentially the same accuracy for $\rho=0.49$, the prefractal BEM being more accurate for $\rho=1/3$, and the Hausdorff BEM being more accurate for $\rho=0.1$. To investigate this further we carried out similar experiments for more values of $\rho$, plotting the ratio of the errors obtained (measured in the $L_{\infty}$ and $L_{2}$ norms) with $\ell=10$ in Figure 16. We observe that:
@@ 6.5 Non-homogeneous or non-disjoint IFS attractors
Figure 17: The scattered field induced by a plane wave, with wavenumber $k=50$ and direction vector $\vartheta=(0,1,-1)/\sqrt{2}$, incident on the “non-homogenous dust” of [25, Fig. 8b, eq. (70)], plotted on three faces of a cube, computed with mesh parameter $h=h_{\mathrm{ref}}=\sqrt{2}/16384$. The scatterer, which is a subset of the plane $\mathbb{R}^{2}\times\{0\}$, is shown in black. Further details are given in §6.5.
Figure 18: $H^{-1/2}_{\Gamma}$ norms of the errors for the two IFS attractors described in §6.5. Top left: $\|\phi_{h}-\phi_{h_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}/\|\phi_{h_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}$, bottom left: $\|\phi_{\ell}-\phi_{\ell_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}/\|\phi_{\ell_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}$, top right: $\|\phi_{h}-\phi_{h/4}\|_{H^{-1/2}_{\Gamma}}$ (see Footnote 18), bottom right: $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}$.
@@
    the “non-homogeneous dust”, shown in Figure 17, and defined in [25, Fig. 8b, eq. (70)] (with $M=4$, $\rho_{1}=\rho_{2}=\rho_{3}=\frac{1}{4}$, $\rho_{4}=\frac{1}{2}$ and Hausdorff dimension $d\approx 1.20$);
@@
    the Sierpinski triangle [16, §6.3] (with $M=3$, $s_{1}(x)=\frac{1}{2}x$, $s_{2}(x)=\frac{1}{2}x+(\frac{1}{2},0)$, $s_{3}(x)=\frac{1}{2}x+(\frac{1}{4},\frac{\sqrt{3}}{4})$ and Hausdorff dimension $d=\log 3/\,\log 2\approx 1.58$);
Both satisfy the OSC (5) and are hence $d$-sets for the stated values of $d$. But only (ii) is homogeneous, and only (i) is disjoint and satisfies the assumptions of the Hausdorff-BEM convergence theory of §5.2. Figure 18 shows the $H^{-1/2}_{\Gamma}$ norms of the Hausdorff-BEM errors computed against a fixed reference solution (left column) and the norms of the differences between Hausdorff-BEM solutions on consecutive meshes as in Figure 11 (right column), for a range of wavenumbers. [^18] In §6.2 we motivated the study of the errors $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}$ between consecutive levels with the bound (129). In the case of non-homogeneous IFSs, denoting by $\phi_{h}$ the Hausdorff-BEM solution corresponding to a mesh of size $h$, assuming that the theoretical error bound (80) is sharp, i.e. $\|\phi-\phi_{h}\|_{H^{-1/2}_{\Gamma}}=Ch^{\alpha}$ for some $C>0$ and $\alpha=(1-n+d)/2$, it follows that $Ch^{\alpha}\big(1-(h^{\prime}/h)^{\alpha}\big)\leq\|\phi_{h}-\phi_{h^{\prime}}\|_{H^{-1/2}_{\Gamma}}\leq Ch^{\alpha}\big(1+(h^{\prime}/h)^{\alpha}\big)$ for $0<h^{\prime}<h$. In the example (i), two consecutive meshes have $h^{\prime}/h=1/4$; thus we expect theoretically that $\|\phi_{h}-\phi_{h^{\prime}}\|_{H^{-1/2}_{\Gamma}}$ is approximately proportional to $h^{\alpha}$. For the Sierpinski triangle (ii) we use meshes of $N=3^{\ell}$ elements of diameter $2^{-\ell}$ and we plot the errors against the level $\ell$ for $\ell=0,\ldots,9$, using a reference solution with $\ell_{\mathrm{ref}}=10$, i.e. $N_{\mathrm{ref}}=3^{10}=59049$. For (i), since the IFS is non-homogeneous the BEM mesh is not parametrised by the level $\ell$ but rather by the mesh size $h=\frac{\sqrt{2}}{4^{j}}$, $j=0,\ldots,5$, giving the number of degrees of freedom as $N=1,7,40,217,1159,6160,$ respectively. The reference solution has $h=h_{\mathrm{ref}}=\frac{\sqrt{2}}{4^{6}}=\frac{\sqrt{2}}{4096}$ and $N_{\mathrm{ref}}=75316$. For quadrature, for (i) we use $h_{Q}=\frac{h}{16}$ and for (ii) $h_{Q}=\frac{h}{4}$.

### v2

@@ 6 Numerical results
In this section we present numerical results that illustrate our methods and assess the sharpness of our theoretical predictions. The Hausdorff BEM (55) has been implemented in the Julia language [5] and the code is available at https://github.com/AndrewGibbs/IFSintegrals. In all examples described in this section, we consider the scattering of plane waves, i.e. the datum $g$ is as in (29) with $u^{i}(x)=e^{{\mathrm{i}}k\vartheta\cdot x}$ and $|\vartheta|=1$. We validate our implementation against a different method in §6.4, but in the rest of our experiments we use as a reference solution a more accurate Hausdorff-BEM solution with a large number of degrees of freedom $N_{\mathrm{ref}}$. Most of our experiments are for homogeneous attractors, in which case $N_{\mathrm{ref}}=M^{\ell_{\mathrm{ref}}}$ for some $\ell_{\mathrm{ref}}\in\mathbb{N}$. For Cantor sets $(M=2)$ we choose $\ell_{\mathrm{ref}}=15$, so that $N_{\mathrm{ref}}=32768$. For Cantor Dusts ($M=4$) we choose $\ell_{\mathrm{ref}}=8$, so that $N_{\mathrm{ref}}=65536$. Since we do not use any matrix compression, the memory required to store the Galerkin matrices grows like $N^{2}$, with $N=M^{\ell}$ when the attractor is homogeneous. Thus, while many of our experiments were run on a standard laptop, some (in particular, the calculation of reference solutions) required the use of the Myriad High Performance Computing Facility available at University College London, which has computing nodes available with 1.5TB of RAM.
@@
Our implementation uses the quadrature rules described in §5.4. Precisely, we approximate the right-hand side in the linear system and linear functionals of the solution (the scattered field and far field) using the quadrature rules (97) and (98), respectively. We approximate the diagonal matrix elements $A_{ii}$ by (93)-(96). To approximate $A_{ij}$ with $i\neq j$ we use (92) with $h_{Q}$ replaced by $h_{Q,i,j}\geq h_{Q}$ given (in terms of $h_{Q}$) by (121)-(123). All these quadrature rules depend on the parameter $h_{Q}$. We choose $h_{Q}=C_{Q}h$, where $C_{Q}>0$ is a constant independent of $h$. While the requirement $h_{Q}=C_{Q}h$ is weaker (i.e. it requires fewer quadrature points) than the conditions $h_{Q}\leq C_{Q}h^{1+d/2}$ (see (111)) and $h_{Q}\leq C_{Q}h^{1+d}$ (see (109)) required by our theory, our numerical experiments suggest that, in practice, $h_{Q}=C_{Q}h$ is sufficient to achieve our theoretical convergence rates, even when using the reduced quadrature of Remark 5.19. Except where indicated otherwise, in the simulations reported below for homogeneous attractors we use $C_{Q}=\rho^{2}$ for cases with $k<20$, $C_{Q}=\rho^{4}$ when $20\leq k\leq 50$.
@@
We measure the accuracy of our BEM solutions in the $H^{-1/2}_{\Gamma}$ norm. These $H^{-1/2}_{\Gamma}$ norms are computed by expressing them in terms of a single-layer BIO with wavenumber $k={\mathrm{i}}$, which we denote $S^{\mathrm{i}}$, as in [19, Table 1]. Practically, we achieve this by assembling an approximation $A^{{\mathrm{i}},Q}$ to the Galerkin matrix $A^{\mathrm{i}}$ for the operator $S^{\mathrm{i}}$, analogous to (91), with $k={\mathrm{i}}$ and $N_{\mathrm{ref}}$ degrees of freedom, using quadrature approximations analogous to (93)-(96) and (92), choosing $h_{Q}$ the same as for the reference solution, with the reduced quadrature formulae (121)-(123) applied as if $k=1$. (While the quadrature convergence analysis in [30, §5] was presented only for $k>0$, one can check that the relevant results also extend, mutatis mutandis, to the case $k={\mathrm{i}}$.) Next, we view $\phi_{N}^{Q}$ as an element of the larger space $V_{N_{\mathrm{ref}}}$ and define $\vec{v}$} as the coefficient vector of $\phi_{N}^{Q}-\phi_{N_{\mathrm{ref}}}^{Q}$ in $V_{N_{\mathrm{ref}}}$. Then, arguing as in [19, Table 1] (and see Footnote 15), it follows that
@@ 6.1 $H^{-1/2}_{\Gamma}$-norm convergence, Cantor set ($n=1$)
In Figure 7 we show an example of the total and scattered fields for the middle-third case ($\rho=1/3$). In Figure 8 we present plots of the relative errors in the Hausdorff-BEM solution for $\ell=0,1,\ldots,14$, measured in $H^{-1/2}_{\Gamma}$ norm, using $\phi_{15}$ as reference solution, for various $\rho$ and $k$ values. In all cases the incident wave is the plane wave $u^{i}(x)={\mathrm{e}}^{{\mathrm{i}}k\vartheta\cdot x}$ with $\vartheta=(1,-1)/\sqrt{2}$. The experiments in Figures 7 and 8 are carried out, as indicated in the second paragraph of this section, using the quadrature rules described in §5.4, with the reduced quadrature of Remark 5.19, and with $h_{Q}=\rho^{2}h$ for $k=0.1$ and $h_{Q}=\rho^{4}h$ for $k=50$. This means that the off-diagonal entries of the BEM matrix require up to $16$ and $256$ evaluations of the Helmholtz fundamental solution for $k=0.1$ and $k=50$, respectively.
@@ 6.2 $H^{-1/2}_{\Gamma}$-norm convergence, Cantor dust ($n=2$)
a $d$-set with $d={\rm dim_{H}}(\Gamma)=\log{4}/\log{(1/\rho)}$, see Figure 1. Such a screen generates a non-zero scattered field if and only if $\rho>1/4$ (see [15, Example 8.2] and the discussion before Lemma 4.4). In Figure 9 we present results similar to those in Figure 8, albeit for a more restricted range of $\ell$ and a lower value of $k$ in the right-hand panel, due to the increased computational cost associated with the change from $M=2$ (Cantor set) to $M=4$ (Cantor dust). The Hausdorff-BEM solution $\phi_{\ell}$ corresponds to a mesh of $N=4^{\ell}$ elements of diameter $h=\sqrt{2}\rho^{\ell}$. The incident wave is the plane wave $u^{i}(x)={\mathrm{e}}^{{\mathrm{i}}k\vartheta\cdot x}$ with $\vartheta=(0,1,-1)/\sqrt{2}$. As indicated at the beginning of this section, we use the quadrature rules described in §5.4, with the reduced quadrature of Remark 5.19, and with $h_{Q}=\rho^{2}h$ for both $k=0.1$ and $k=5$. As a result, the off-diagonal entries of the BEM matrix require up to $256$ evaluations of $\Phi$.
Figure 9: $H^{-1/2}_{\Gamma}$ relative errors for Hausdorff-BEM solutions on Cantor dusts.
@@
$\|\phi-\phi_{\ell}\|_{H^{-1/2}_{\Gamma}}=C\mathfrak{a}^{\ell},\quad\ell\in\mathbb{N},$ (126)
for $\mathfrak{a}:=(4\rho)^{-1/2}\in(1/\sqrt{2},1)$ and some $C>0$. Then, by the triangle inequality,
$C\mathfrak{a}^{\ell}(1-\mathfrak{a}^{\ell^{\prime}-\ell})\leq\|\phi_{\ell}-\phi_{\ell^{\prime}}\|_{H^{-1/2}_{\Gamma}}\leq C\mathfrak{a}^{\ell}(1+\mathfrak{a}^{\ell^{\prime}-\ell}),\qquad\ell,\ell^{\prime}\in\mathbb{N}_{0}\;\text{with}\;\ell<\ell^{\prime},$ (127)
@@
In Figure 11 we plot, as a better test of the sharpness of our theory, the norm $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}$, for $\ell=0,\ldots,7$, of the difference between Galerkin solutions at consecutive levels $\ell$. We observe, for each choice of $\rho$ and $k$, that, for large enough $\ell$ (how large depending on $\rho$ and $k$), $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}=c\mathfrak{a}^{\ell}=c(4\rho)^{-\ell/2}$, for some constant $c>0$, also depending on $\rho$ and $k$, in agreement with (127) with $\ell^{\prime}=\ell+1$. Moreover, since $\|\phi-\phi_{\ell}\|_{H^{-1/2}_{\Gamma}}\to 0$ as $\ell\to\infty$ by Theorem 5.6, this observed convergence implies that
$\|\phi-\phi_{\ell}\|_{H^{-1/2}_{\Gamma}}\leq\sum_{m=\ell}^{\infty}\|\phi_{m}-\phi_{m+1}\|_{H^{-1/2}_{\Gamma}}=\frac{c}{1-\mathfrak{a}}\,\mathfrak{a}^{\ell},$
@@ 6.3 Convergence of the scattered field
In Figure 13 we show $L_{\infty}$ errors in the near- and far-field for scattering by Cantor sets and Cantor dusts, obtained by computing the maximum error over a suitable set of sample points. In more detail, define the parameter $\mathcal{N}(k):=10\max(k,2)$, which is always an even integer for the values of $k$ in our experiments. For the near-field, when $n=1$ we sample at $4\mathcal{N}$ points on the boundary of the square $(-1,2)\times(-1.5,1.5)$, and when $n=2$ we sample at $\mathcal{N}^{2}$ points on a uniform grid on the square $(-1,2)\times(-1,2)\times\{-1\}$ (recall that $\Gamma\subset[0,1]\times[0,1]\times\{0\}$). For the far-field, when $n=1$ we sample at $\mathcal{N}$ points on the circle $\mathds{S}^{1}$, and when $n=2$ we sample at $\frac{\mathcal{N}}{2}\times\mathcal{N}$ points on the sphere $\mathds{S}^{2}$, chosen such that the points form a uniform grid in spherical coordinate space $0,\pi]\times[0,2\pi]$.
@@ 6.4 Comparison against “prefractal BEM”
Figure 16: Ratio of errors of Hausdorff-BEM (${u}_{10}^{\infty}$) and prefractal-BEM ($\tilde{u}_{10}^{\infty}$) approximations of the far-field pattern $u^{\infty}$ with $\ell=10$ for scattering by Cantor sets with $k=0.1$, $k=30$ and $k=50$ and $\vartheta=(1/2,-\sqrt{3}/2)$, for a range of values of $\rho\in(0,1/2)$, measured in the $L_{\infty}$ (left panel) and $L_{2}$ (right panel) norms on $\mathds{S}^{1}$. A value below 11 indicates that the Hausdorff BEM is more accurate, while a value above 11 indicates that the prefractal BEM is more accurate. In both cases $\ell_{\mathrm{ref}}=13$.
@@
In Figure 15 we show $L_{\infty}$ errors in the far-field pattern between the two methods for $\rho=0.1$, $1/3$ and $0.49$, with incident direction $\vartheta=(1/2,-\sqrt{3}/2)$ and wavenumbers $k=0.1$ and $k=50$ (with $h_{Q}=\rho^{6}h$ for $k=0.1$ and $h_{Q}=\rho^{8}h$ for $k=50$ for the Hausdorff BEM), obtained by computing the maximum error over 300 equally-spaced observation angles in $\mathds{S}^{1}$. For each method we calculate the errors in two different ways, using, firstly, a Hausdorff-BEM reference solution, and, secondly, a prefractal-BEM reference solution; in both cases the reference solution is computed with $\ell_{\mathrm{ref}}=13$. In all cases, it is clear that both methods are converging to the same solution, validating both methods. Indeed, it appears that both methods converge at the same $2^{-\ell}$ rate. However, the results suggest that the relative accuracy of the two methods is dependent on the value of $\rho$, with the two methods having essentially the same accuracy for $\rho=0.49$, the prefractal BEM being more accurate for $\rho=1/3$, and the Hausdorff BEM being more accurate for $\rho=0.1$. To investigate this further we carried out similar experiments for more values of $\rho$, plotting the ratio of the errors obtained (measured in the $L_{\infty}$ and $L_{2}$ norms) with $\ell=10$ in Figure 16. We observe that:
@@ 6.5 Non-homogeneous or non-disjoint IFS attractors
Figure 17: The scattered field induced by a plane wave, with wavenumber $k=50$ and direction vector $\vartheta=(0,1,-1)/\sqrt{2}$, incident on the “non-homogenous dust” of [30, Fig. 8b, eq. (70)], plotted on three faces of a cube, computed with mesh parameter $h=h_{\mathrm{ref}}=\sqrt{2}/16384$. The scatterer, which is a subset of the plane $\mathbb{R}^{2}\times\{0\}$, is shown in black. Further details are given in §6.5.
Figure 18: $H^{-1/2}_{\Gamma}$ norms of the errors for the two IFS attractors described in §6.5. Top left: $\|\phi_{h}-\phi_{h_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}/\|\phi_{h_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}$, bottom left: $\|\phi_{\ell}-\phi_{\ell_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}/\|\phi_{\ell_{\mathrm{ref}}}\|_{H^{-1/2}_{\Gamma}}$, top right: $\|\phi_{h}-\phi_{h/4}\|_{H^{-1/2}_{\Gamma}}$ (see Footnote 17), bottom right: $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}$.
@@
    the “non-homogeneous dust”, shown in Figure 17, and defined in [30, Fig. 8b, eq. (70)] (with $M=4$, $\rho_{1}=\rho_{2}=\rho_{3}=\frac{1}{4}$, $\rho_{4}=\frac{1}{2}$ and Hausdorff dimension $d\approx 1.20$);
@@
    the Sierpinski triangle [19, §6.3] (with $M=3$, $s_{1}(x)=\frac{1}{2}x$, $s_{2}(x)=\frac{1}{2}x+(\frac{1}{2},0)$, $s_{3}(x)=\frac{1}{2}x+(\frac{1}{4},\frac{\sqrt{3}}{4})$ and Hausdorff dimension $d=\log 3/\,\log 2\approx 1.58$);
Both satisfy the OSC (4) and are hence $d$-sets for the stated values of $d$. But only (ii) is homogeneous, and only (i) is disjoint and satisfies the assumptions of the Hausdorff-BEM convergence theory of §5.2. Figure 18 shows the $H^{-1/2}_{\Gamma}$ norms of the Hausdorff-BEM errors computed against a fixed reference solution (left column) and the norms of the differences between Hausdorff-BEM solutions on consecutive meshes as in Figure 11 (right column), for a range of wavenumbers. [^17] In §6.2 we motivated the study of the errors $\|\phi_{\ell}-\phi_{\ell+1}\|_{H^{-1/2}_{\Gamma}}$ between consecutive levels with the bound (127). In the case of non-homogeneous IFSs, denoting by $\phi_{h}$ the Hausdorff-BEM solution corresponding to a mesh of size $h$, assuming that the theoretical error bound (78) is sharp, i.e. $\|\phi-\phi_{h}\|_{H^{-1/2}_{\Gamma}}=Ch^{\mathfrak{a}}$ for some $C>0$ and $\mathfrak{a}=(1-n+d)/2$, it follows that $Ch^{\mathfrak{a}}\big(1-(h^{\prime}/h)^{\mathfrak{a}}\big)\leq\|\phi_{h}-\phi_{h^{\prime}}\|_{H^{-1/2}_{\Gamma}}\leq Ch^{\mathfrak{a}}\big(1+(h^{\prime}/h)^{\mathfrak{a}}\big)$ for $0<h^{\prime}<h$. In the example (i), two consecutive meshes have $h^{\prime}/h=1/4$; thus we expect theoretically that $\|\phi_{h}-\phi_{h^{\prime}}\|_{H^{-1/2}_{\Gamma}}$ is approximately proportional to $h^{\mathfrak{a}}$. For the Sierpinski triangle (ii) we use meshes of $N=3^{\ell}$ elements of diameter $2^{-\ell}$ and we plot the errors against the level $\ell$ for $\ell=0,\ldots,9$, using a reference solution with $\ell_{\mathrm{ref}}=10$, i.e. $N_{\mathrm{ref}}=3^{10}=59049$. For (i), since the IFS is non-homogeneous the BEM mesh is not parametrised by the level $\ell$ but rather by the mesh size $h=\frac{\sqrt{2}}{4^{j}}$, $j=0,\ldots,5$, giving the number of degrees of freedom as $N=1,7,40,217,1159,6160,$ respectively. The reference solution has $h=h_{\mathrm{ref}}=\frac{\sqrt{2}}{4^{6}}=\frac{\sqrt{2}}{4096}$ and $N_{\mathrm{ref}}=75316$. For quadrature, for (i) we use $h_{Q}=\frac{h}{16}$ and for (ii) $h_{Q}=\frac{h}{4}$.